How many milliliters of 9.26 M nitric acid solution should be used to prepare 5.00 L of 0.600 M ?


1
Expert's answer
2021-09-17T02:12:49-0400

C1*V1=C2*V2


C1=9.6 M

V2=5.00 L

C2=0.6 M

V-?

V1=C2*V2/C1

=0.6*5/9.6=0.32397 L


If 1 L= 1000 mL

0.32397 L= ?

(0.32397*1000)

=323.97 mL

Answer: 323.97 mL 9.26 of M nitric acid solution should be used